OpenStudy (anonymous):

If a buffer solution is 0.270 M in a weak acid (Ka=6.2 x 10-5) and 0.540 M in its cojugate acid, what is the pH?

OpenStudy (anonymous):

pH=6.2*10^5+log([.250]\[.540])= -.33

OpenStudy (jfraser):

flip the fraction upside-down, it should be \[pH = pKa + \log(\frac{A^-}{HA})\]